jigglypuff314 (jigglypuff314):

I've just started learning about integrals... so what does it mean by: "Solve the differential equation y' = 4x^(-1/2) at point (1, 12)" ?

OpenStudy (abb0t):

What it means is that you have: \(\sf \frac{dy}{dx}=4x^{-\frac{1}{2}} \) this is called a seperable differential equation.

OpenStudy (abb0t):

(1,12) is your initial conditions.

jigglypuff314 (jigglypuff314):

so I find the anti derivative and plug in (1,12) to find C ?

OpenStudy (abb0t):

I mean, you can think of it as "cross-multiplying" but you shouldn't be thinking like that now at differential equations level. Lol \[\sf dy=(4x^{-\frac{1}{2}})~dx\]

OpenStudy (abb0t):

You separated them into their respective variable. Now, how do you solve for any antiderivative? You integrate BOTH SIDES

jigglypuff314 (jigglypuff314):

I can get that y = 8x^(1/2) + C I was just confused at the terminology of What is the question asking for?

OpenStudy (abb0t):

\[\sf \int dy = \int (4x^{-\frac{1}{2}})dy\]

OpenStudy (abb0t):

Sorry, the second variable should be \(dx\) not \(\color{red}{dy}\)

ganeshie8 (ganeshie8):

more generally the question is begging u to find a function \(y\) with below properties: 1) the function \(y\) has a slope \(4x^{-1/2}\) 2) function passes thru the point \( (1,2)\)

jigglypuff314 (jigglypuff314):

so would the equation be y = 8x^(1/2) + 4 ?

OpenStudy (abb0t):

That is a relatively simple integral which you should be able to solve. Then, plug in the initial conditions y = 12, x = 1 and solve for C.